# Do We Need All the Synthetic Data? Towards Targeted Synthetic Image Augmentation via Diffusion Models

### v1

@@ F.2 Additional experimental results
Effect of the upsampling factors. Table 2 illustrates the performance of our method and Useful when varying the upsampling factors. As discussed in Section 3.2, Useful achieves the best performance at $k=2$ due to overfitting noise at larger $k$. On the other hand, using synthetic images, our method benefits from larger values of $k$, yielding the best performance at $k=5$ for the CIFAR datasets and $k=4$ for the Tiny-ImageNet dataset. This empirical result reinforces our theoretical findings in Section 4.2.
@@
k Cifar10 (Useful) Cifar10 (Ours) Cifar100 (Ours) Tiny-ImageNet (Ours)
@@
Figure 5: Variance of mini-batch gradients of ResNet18 with SGD on the augmented CIFAR10 dataset found by Useful and Ours when varying the upsampling factor $k$
@@
Variance of mini-batch gradients. Figure 5 presents the mini-batch gradient variance during training of ResNet18 using SGD on the augmented CIFAR-10 datasets selected by Useful and our proposed method. Across all upsampling factors $k$, our method consistently yields lower gradient variance throughout the entire training process. This indicates that our synthetic data provides a more stable and consistent learning signal, which can contribute to more effective optimization. Notably, the variance gap between our method and Useful widens significantly after epochs 100 and 150—key points at which the learning rate is decayed. This suggests that our method enables the model to adapt more smoothly to changes in the learning rate, likely due to the benign noise introduced in the faithful synthetic images. Lower gradient variance is often associated with more stable convergence and improved generalization, highlighting the advantage of our approach.

### v2

@@ F.2 Additional experimental results
State-of-the-art architectures. To further demonstrate the effectiveness of our method, we evaluated it on two widely used modern vision backbones: ConvNeXt-T (Liu et al., 2022), a convolutional network that revisits ConvNet design with architectural refinements inspired by transformers, and Swin-T (Liu et al., 2021), a hierarchical vision transformer that introduces shifted windows for efficient and scalable self-attention. Both models were trained on CIFAR-10 using SGD with $k=2$. As shown in Table 2, our method consistently outperforms both the Original (no augmentation) and Upsample. In particular, for ConvNeXt-T, our approach achieves a nearly 7% reduction in test error compared to the second-best method, highlighting its strong compatibility with modern architectures.
Qualitative results. Figure 6 presents examples of slow-learnable (top) and fast-learnable (bottom) samples in CIFAR-10 where the slow-learnable samples are specifically targeted in our synthetic data augmentation. Slow-learnable samples are often visually ambiguous: the object may be partially visible, relatively small compared to the background, or blended with clutter, making them harder to recognize. In contrast, fast-learnable samples are clear and representative of their class, with the object occupying most of the image and little background interference. These characteristics explain why slow-learnable samples are acquired later in training, whereas fast-learnable ones are learned quickly. Figure 6 shows that our synthetic examples preserve key visual content (features)—such as object shape, structure, and spatial arrangement while introducing subtle variations in texture or color (noise). For instance, an image of a bird in the second row is reproduced with the same layout and pose, but with different color for the neck part. Such nuanced transformations are difficult to achieve with standard augmentations like random cropping or flipping, highlighting the value of generative augmentation in our approach.
Effect of the upsampling factors. Table 5 illustrates the performance of our method and Upsample when varying the upsampling factors. As discussed in Section 4, Upsample achieves the best performance at $k=2$ due to overfitting noise at larger $k$. On the other hand, using synthetic images, our method benefits from larger values of $k$, yielding the best performance at $k=5$ for the CIFAR datasets and $k=4$ for the Tiny-ImageNet dataset. This empirical result reinforces our theoretical findings in Section 4.2.
@@
k Cifar10 (Upsample) Cifar10 (Ours) Cifar100 (Ours) Tiny-ImageNet (Ours)
@@
Table 7: Training ResNet18 on CIFAR10 with different data selection strategy. We used SGD and set the upsampling factor $k$ to 5 here.
Method Misclassification High loss Slow-learnable (Upsample and Ours)
Test error 4.87 $\pm$ 0.10 4.77 $\pm$ 0.06 4.42 $\pm$ 0.04
Alternative selection strategies. Our approach differs from core-set selection methods Guo et al. (2022), which aim to reduce dataset size while matching the performance of training on the full data. Such methods either require training the model (or a smaller proxy) and use its statistics to find the coreset or update the coreset during the training. In contrast to coreset selection, the goal of our approach is to improve the generalization performance over that of full data, by reducing the simplicity bias of training. We showed that this can be done by augmenting the data with faithful synthetic examples corresponding to the slow-learnable part of the data. Motivated by theory, we found such examples by clustering the model output early in training (without training a proxy model or updating the subset during training). However, other more heuristic approaches can be also used to find slow-learnable examples. In this section, we conducted new experiments to find slow-learnable examples based on (1) high loss and (2) misclassification, at the same checkpoint as used in our experiments in the paper. As shown Table 7, our method outperforms both heuristic-based selection strategies. Corset selection methods can be applied on top of our synthetically augmented data to further improve data-efficiency. This is an exciting direction we leave for future work.
Figure 7: Variance of mini-batch gradients of ResNet18 with SGD on the augmented CIFAR10 dataset found by Upsample and Ours when varying the upsampling factor $k$. Our method yields lower gradient variance throughout the entire training process.
@@
Variance of mini-batch gradients. Figure 7 presents the mini-batch gradient variance during training of ResNet18 using SGD on the augmented CIFAR-10 datasets selected by Upsample and our proposed method. Across all upsampling factors $k$, our method consistently yields lower gradient variance throughout the entire training process. This indicates that our synthetic data provides a more stable and consistent learning signal, which can contribute to more effective optimization. Notably, the variance gap between our method and Upsample widens significantly after epochs 100 and 150—key points at which the learning rate is decayed. This suggests that our method enables the model to adapt more smoothly to changes in the learning rate, likely due to the benign noise introduced in the faithful synthetic images. Lower gradient variance is often associated with more stable convergence and improved generalization, highlighting the advantage of our approach.